OpenStudy (anonymous):

somebody please help me.. i really need it bad Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (imstuck):

First find the slope of the line using the slope formula. then use that slope to write the equation from the points. The slope is\[\frac{ 1-2 }{ 2-(-3) }=-\frac{ 1 }{ 5 }\]Use that now in your point-slope formula\[y-y _{1}=m(x-x _{1})\]

OpenStudy (imstuck):

m is the slope you found of -1/5. Use one of the points for your y1 and x1. Let's pick (2,1) because it doesn't have any negatives in it.\[y-1=-\frac{ 1 }{ 5 }(x-2)\]Oh well, they used the other point, so let's do that one, too.

OpenStudy (imstuck):

\[y-2=-\frac{ 1 }{ 5 }(x+3)\]

OpenStudy (imstuck):

It's the last choice of your options.

OpenStudy (anonymous):

ok thankyou so much